OpenStudy (anonymous):

Please Help! solve: 3tan(x/2)+3=0 for 0 < theta <180degrees

OpenStudy (anonymous):

3tan(x/2)+3=0 3 tan(x/2)=-3 tan(x/2)=-1 Now, let y=x/2 then, tan(y)=-1 , 0<y<90

OpenStudy (anonymous):

Now, solve for y can u?

OpenStudy (anonymous):

think i am confused on how to go about solving for y..

OpenStudy (anonymous):

if you take the inverse of tan to both side you will get y =-45 tan(y)=-1 arctan(y)=arctan(-1) y=-45 and since 0<y<90 then y=180-45=135

OpenStudy (anonymous):

which are out of range So, NO SOLUTION of y.. Thus, NO solution of x

OpenStudy (anonymous):

this range 0<y<90 for y or x in the origin problem 0<y<90 or 0<x<90

OpenStudy (anonymous):

Since, 0<x<180 and y=x/2 range of y is 0<y<90